Circular No. 6743

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

POSSIBLE NEW SOFT γ-RAY REPEATER K. Hurley, Space Sciences Laboratory, Berkeley, on behalf of the Ulys-

ses Gamma-Ray Burst Team; C. Kouveliotou, Universities Space Research Association (USRA), on behalf of the Gamma-Ray Observatory BATSE team; and T. Cline and E. Mazets, on behalf of the KONUS-WIND team, report: “Ulysses has observed three similar short intense bursts whose triangulated positions are consistent with a single location. The first two occurred on June 29 (and were also observed by KONUS-WIND), and the third on Sept. 12; the second and third are the BATSE triggers described below. For each burst, we can derive a single annulus of possible arrival directions. They intersect over a region defined by the following coordinates (equinox 2000.0): α = 18^{h}14^{m}50^{s}, δ = −13^{o}36′.6; 18^{h}14^{m}25^{s}, −14^{o}01′.7; 18^{h}15^{m}26^{s}, −13^{o}15′.0; 18^{h}15^{m}02^{s}, −13^{o}39′.8.”

C. Kouveliotou, USRA and Marshall Space Flight Center (MSFC), NASA; G. J. Fishman and C. A. Meegan, MSFC; and P. Woods, University of Alabama in Huntsville, report on behalf of the BATSE Team: “BATSE triggered on two very intense, short, soft events on June 29.27189 and Sept. 12.25652 UT. Each lasted ∼ 2 s, and their (preliminary) spectra can be fitted with a power-law with indices of ∼ −3.8 and −4.2, respectively. Given the characteristics and common location noted above, we classify them as soft-γ-ray-repeater events from a previously unknown source.”

D. A. Smith, A. M. Levine, E. H. Morgan, and R. A. Remillard, Massachusetts Institute of Technology (MIT), on behalf of the RXTE/ASM team at MIT and Goddard Space Flight Center, NASA; and R. Rutledge, Max-Planck-Institut für Extraterrestrische Physik, write: “The RXTE All Sky Monitor has detected the Sept. 12 event reported above; ASM camera 3 detected a bright x-ray flash that lasted ∼ 2.8 s and yielded a peak countrate of ∼ 270 counts/s. The inferred peak flux, including corrections for background and location in the field-of-view, is ∼ 10 Crab (2–12 keV). We obtain a single long, narrow line of position. Our formal analysis gives an error box with corners at α = 18^{h}25^{m}39^{s}.0, δ = −14^{o}51′48′′; 18^{h}25^{m}42^{s}.6, −14^{o}48′50′′; 18^{h}34^{m}24^{s}.9, −15^{o}28′10′′; 18^{h}34^{m}04^{s}.5, −15^{o}23′31′′ (equinox 2000.0). However, the determination of the ends of the error box along the long direction is subject to systematic errors of ∼ 2^{o}. With this consideration, the extension of our error box by ∼ 2^{o}.8 in the long direction is consistent with the intersection of the position annuli from the IPN reported above. The combined IPN/ASM error box is centered at α = 18^{h}14^{m}.7, δ = −13^{o}40′ and has a radius (90-percent confidence) of 5′.”

1997 September 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green